02 Feb 2023 2nd half ECD

Night center: 00:33 MST

Morning 12-deg twilight: 06:18 MST

Sunrise: 07:10 MST

Summary: DX adSec was used during the first half of the night (LBTI) and the start of the ECD time, but around 10:00 UT there was a fast diagnostic failure (IT 8825), after which DX AO operations were halted and only SX-LUCI1-AO tasks were done: SX-SOUL optimize gain tuning. TCS patches were activated (PCS and OSS) and the AO software changes made. The TCS and AO updates will all remain in place. See 20230202 UT Second-half ECD Night Log for details.

Tasks to be done:

Google doc spreadsheet with outline of plan

DX-SOUL (recoup lost Comm time; 6 hrs - a full 1/2 night)

  • AO conditions seeing < 1.2", cirrus
  • SOUL-LUCI_Comm_OT_V#.xml
  • binocular presets, can work on( ✅ SX optimize gain tuning ) in parallel with DX work
  • depends on DX adSec condition (IT 8825)
  • Feasibility needs to be verified with the Arcetri team. This would be top priority.

✅ SX-SOUL

  • ✅Optimize Gain Optimization Tuning
  • ✅OSCO

Verify optimize gain tuning

  • LUCI1-AO

DX: Full AO Field Aberration Check (0.5hr)

Jenny wrote that Enrico mentioned that DX has always had this "speckle pattern" although this is not something we recall seeing. We are unsure if this is the result of contamination from the window or of some other optic in LUCI (detector focus is at limits for N30). We would like to execute a full AO bin1 test with dithers around the full N30 FoV to investigate the IQ across the field.
  • Scripts pending generation, but likely M61 for fun.
  • Seeing <1.2"

ESM testing (3 hrs)

  • Testing using the ESM_V57.xml OT, testing regular ESM optimization and data quality
  • Targets M61, M67
  • Can be done in parallel with AO SX test
  • Seeing <2"

All-modes ESM (3 hrs)

  • LUCI1-ESM & LUCI2-SL
  • ESM_V57.xml M92 or M67

Check edges of LUCI-AGw patrol fields (3.5 hrs)

  • SL: both LUCI1 and LUCI2
  • poor weather
  • target prepared: NGC3190 @ RA=10, DEC=+21

[SCRATCH; bad GPS (IT8819)] Pointing Models (telescope needs to be in good thermal equilibrium)

  • collect data for a new LUCI/MODS pointing model
  • Prerequisite: GPS has to be restored, maybe Friday...
  • LUCIs
    • N375 cameras, N30_FieldStop, K-band, DIT/NDIT = 2.51s/2, FlexComp off
    • Before starting, measure rotator center using at least two track mode presets at diametrically opposed PAs
    • collimate, then collect data - provides collimation and pointing model.

Photometric standards (requires photometric conditions)

TMS

  • Prerequisite: Need to check alignment post power-outage
  • Channel-dropping on-sky with TMS updates mid-exposure:
    • selectively drop and bring back RR pairs and monitor the effect of the different configurations
    • tms_loop.py -o (-o is needed to ignore the shutter state and send corrections during exposures).
  • Thermal aberrations (need to analyze current data first)

Software release

✅ TCS-related

There are several TCS patches & updates, as listed below. The plan is to release the 3 patches and do items 4,5 and 6 on Wednesday morning, after which the TCS will have to be restarted. (On Feb 1, the GCS patch could not be released). The 3 (now 2) patches (OSS, PCS) will only be activated in the middle of the night, during the handover from LBTI science to ECD time.
  1. OSS patch (improved simulation mode for ARGOS swingarm)
  2. [DEFERRED] GCS patch (addressing IT7752 GCS error messages like "gcs.guidethread.saveImageThread.error left NFS stalled" and "gcs.guidethread.saveImageThread.error right NFS stalled")
  3. PCS patch (addressing IT7380 copointing errors after instrument authorization changes):
    1. copointing errors alarm after instrument authorization changes should be suppressed (it is still being logged but no longer sounds an alarm)
  4. MCS configuration update (minor)
  5. event.dat update (related to PCS patch)
  6. ARGOS white light source proxy update (preserves the power status of the light source in OSS and ARGOS interface)
    1. The ARGOS light source status should now be accurate in OSS (and ARGOS interface).
To activate the patches:

  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-build latest PCS OSS
During the process, it should notify you that it is switching to PCS.20230201 and OSS.20230201. Please answer 'y' to the restart question in step 3) so that the OSS and PCS subsystems will be restarted. On the OSSGUI, if you click Help->About, you should see the OSS version number is now 6.3. There is unfortunately no version number info on the PCSGUI.

In case of problems, you can revert them back (and let the software group know about your issues):

  1. ssh tcs@tcs1.mountain.lbto.org
  2. cd /lbt/tcs/current
  3. ./tcs-build revert PCS OSS
Thanks and have a great day,

✅ AO-related

This link points to the details of AO-related SW updates and how to activate them. These can run in parallel with AO observing. The test version can be switched on before other AO activities, and the ao-supervisor software only needs to be changed on sxadsec.
  1. HK power off the ASM when HK variables time out.
  2. AdSec arbitrator - disable RR mode when it starts up.
  3. Possibly other updates, pending daytime tests.
The versions of the software to be checked out Wednesday night during DEC time are ready on sxadsec, dxadsec, soul-sxwfs, and soul-dxwfs. Details are here: https://wiki.lbto.org/AdaptiveOptics/AOSWTest-20230201-checkout

If there are any problems with the AdSec, let Guido and Xianyu know. If there are any problems with the WFS related to the upgrade, let Brandon know.

On each machine, to switch to the correct version:

adsc_stop # or w_stop_all

use_soul checkout

adsc_start # or w_start_all

At the end of the AO activities, or if anything goes wrong, to revert:

adsc_stop # or w_stop_all

use_soul stable # or "use_soul mixed_mode_control" on dxadsec

adsc_start # or w_start_all

Night Log

Link to google doc log: 20230202 UT Second-half ECD Night Log

-- OlgaKuhn - 31 Jan 2023
Topic revision: r8 - 03 Feb 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback